OpenStudy (anonymous):

5/6x + 1/9 = -1/6x-1/6

hero (hero):

add like terms to the same side producing 5/6x + 1/6x = -1/6 - 1/9

hero (hero):

Actually, even better, just subtract 5/6x from both sides to produce: \[\frac{1}{9} = -\frac{1}{6}x - \frac{1}{6} - \frac{5}{6}x\] then simplify the right side to get \[\frac{1}{9} = -\frac{6x + 1}{6}\] From there, cross multiply.

OpenStudy (anonymous):

thank you